OpenStudy (anonymous):

Determine the values of a and b for which the following function is continuous at every real number x f(x) { ax+b if x ≤ 1 -2bx^2 - 8a if 1 < x < 2 4b^2/a -5x if x ≥ 2 }

OpenStudy (anonymous):

Can someone please help me ? :(

OpenStudy (sirm3d):

\[f(x)=\begin{cases} (1) ax+b & , x\leq 1\\(2) -2bx^2-8a &,1<x<2\\(3) 4b^2/a-5x &, 2\leq x \end{cases}\] let \(x=1\) and set equation (1) = equation (2) let \(x=2\) and set equation (2) = equation (3) solve the two equations.

OpenStudy (anonymous):

@sirm3d

OpenStudy (anonymous):

isn't it a(1)+b = -2a(1)^2 -8a ?

OpenStudy (sirm3d):

my mistake. \[a(1)+b=-2b(1)^2-8a\\a+b=-2b-8a\\3b=-9a\\b=-3a\]

OpenStudy (sirm3d):

equating (2) and (3) when \(x=2\) and \(b=-3a\) \[-2(-3a)(2)^2-8a=\frac{4(-3a)^2}{a}-5(2)\\24a-8a=36a-10\\16a=36a-10\\10=20a\\\frac{1}{2}=a\]

OpenStudy (abb0t):

You could use limits also.

OpenStudy (anonymous):

can you show me how to do it using limits ? @abb0t

OpenStudy (ash2326):

@alyannahere Are you here?

OpenStudy (anonymous):

yes i am @ash2326 pls help me. and most preferably use limits :)

OpenStudy (ash2326):

Yes, it's easy to solve this using limits, \[f(x)=\begin{cases} (1) ax+b & , x\leq 1\\(2) -2bx^2-8a &,1<x<2\\(3) 4b^2/a-5x &, 2\leq x \end{cases}\] For this Let's find the limit at x=1, for f(x) to be continuous the left hand and right hand limit should be equal

OpenStudy (ash2326):

Do you get this part?

OpenStudy (anonymous):

yes. so it's going to be ax+b = a + b x->1^+ ax+b = -a + b x->1^_ right ?

OpenStudy (ash2326):

For left hand limit use ax +b x<=1 For right hand limit use -2bx^2-8a 1<x<2 Both should be equal

OpenStudy (ash2326):

Just put x=1 and equate the two limits

OpenStudy (anonymous):

a=0 ?

OpenStudy (ash2326):

No, you can't find a from this, you need one more equation For that find the limits at x=2 ( the equate the left and right hand limit )

OpenStudy (anonymous):

i'm so sorry but i don't get you @ash2326

OpenStudy (anonymous):

:(

OpenStudy (ash2326):

Limit at x=1 Left hand limit (ax+b), put x=1 a+b Right hand limit -2bx^2-8a put x=1 -2b-8a For the function to be continous \[a+b=-2b-8a\] Do you get this part?

OpenStudy (anonymous):

yup! and that's what i used to solve a but apparently you said it still has to be equal.. which i don't get already

OpenStudy (ash2326):

This is first equation we need one more equation to find a and b Now let's find limit at x=2 Left hand limit -2bx^2-8a put x=2 -8b-8a Right hand limit 4b^2/a-5x put x=2 \[4b^2/a-10\] Now equate the two \[-8b-8a=4b^2/a-10\] Now use this and the previous equation to solve for a and b, Can you solve this now ?

OpenStudy (anonymous):

okay i'll solve it and get back to you @ash2326 :D

OpenStudy (ash2326):

Cool :D

OpenStudy (anonymous):

a= 1/2 b= 3/2 ?

OpenStudy (anonymous):

then after i get these... what do i do now ? ...to get the continuity

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

i am getting A=1/2 AND B=-3/2

OpenStudy (anonymous):

sorry use a and b (small)

OpenStudy (anonymous):

@alyannahere

OpenStudy (anonymous):

oh yeah since it's b=-3a

OpenStudy (anonymous):

would you know how to do continuity @nitz ?

OpenStudy (anonymous):

oops my bad. b=-3/2

OpenStudy (anonymous):

see we have determined the values of a and b assuming that the function is continuous at x=1 and x=2....that is why we have substituted x=1 in first two equations and x=2 in last two......because of continuity \[\lim_{x \rightarrow 1-}f(x)=\lim_{x \rightarrow 1+}f(x)\] ie left and right hand limits are equal

OpenStudy (anonymous):

then ? so it's all equal?at a=1/2 b=-3/2 ... just like that ?

OpenStudy (anonymous):

ya ....you were to find the values of a and b ....and that we have got

OpenStudy (anonymous):

oh thank you ! @nitz

OpenStudy (anonymous):

you are welcome....any other help ..

OpenStudy (anonymous):

yes we are having a text on continuity tom and i don't seem to understand the concept.

OpenStudy (anonymous):

which concept...

OpenStudy (anonymous):

find the largest union of intervals at which the following functions is continuous. Show complete algebraic solution { |x+2| x<=0 3 0<x<=2 squareroot of (3x-15) x>2

OpenStudy (anonymous):

continuity @nitz

OpenStudy (anonymous):

it is discontinuous at x=0 and x=2

OpenStudy (anonymous):

why ?

OpenStudy (anonymous):

discontinuous at x=0 and x=2 because left hand and right hand limits are not equal...

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0-}f(x)=\lim_{x \rightarrow 0-}\left| x+2 \right|=\lim_{h \rightarrow 0}\left| (0-h)+2 \right|=\lim_{h \rightarrow 0}\left| -h+2 \right|=\lim_{h \rightarrow 0}\ \left| 0+2 \right|=2\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0+}f(x)=3\]

OpenStudy (anonymous):

so.left hand limit is 2 and right hand limit is 3 which are not equal and hence the function is discontinuous at x=0

OpenStudy (anonymous):

on a similar basis,....left hand limit at x=2 is 3 and right hand limit is 3(2)-15=-9 ,,,so again both are not equal and function is discontinuous at x=2 also

OpenStudy (anonymous):

right?/

OpenStudy (anonymous):

@alyannahere

OpenStudy (anonymous):

so ....the function is continuous at all points except x=0 and x=2

OpenStudy (anonymous):

R-{0,2}

OpenStudy (anonymous):

are we not going to consider squareroot of (3x-15) x>2 ?

OpenStudy (anonymous):

sorry i forgot the squareroot..

OpenStudy (anonymous):

lol :))

OpenStudy (anonymous):

but still when we check x=2 for continuity ...left hand limit ie for x<=2 the value of function is 3 and for right hand limit sqrt(3x-15)=sqrt(3(2)-15)=sqrt(6-15)=sqrt(-9)=3i....still as 3 is not equal to 3i,....so left and right hand limits are not equal and hence function is discontinuous at x=2

OpenStudy (anonymous):

any doubt

OpenStudy (anonymous):

no i think i understand it already. thanks @nitz !

OpenStudy (anonymous):

you are welcome....for any help u can refer these tutorials http://tutorial.math.lamar.edu/Classes/CalcI/Continuity.aspx